OpenStudy (christos):

WHERE IS MY MISTAKE? :SS 4^x - 2^x - 12 = 0 4^x - 2^x = 12 log(4^x) - log(2^x) = log(12) x*log(4) - x*log(2) = log(12) x*(log(4) - log(2)) = log(12) x*log(2) = log(12) x = log(12) / log(2) x = log_2(12) 2^x = 12 WHERE IS MY MISTAKE TILL NOW?

OpenStudy (anonymous):

line one

OpenStudy (christos):

what is it

OpenStudy (anonymous):

you cannot take the log of both sides by taking the log separately

OpenStudy (anonymous):

in other words if \(x+y=z\) it is not true that \[\log(x)+\log(y)=\log(z)\]

OpenStudy (christos):

what is it was x + y - z = 0 ?

OpenStudy (christos):

what if*

OpenStudy (whpalmer4):

Convert the 4^x and 2^x to be expressions of the same base...

OpenStudy (anonymous):

the gimmick here is to rewrite \[4^x+2^x-12=0\] as \[2^{2x}-x^x-12=0\] and solve the quadratic equation

OpenStudy (christos):

what if it was x + y - z = 0 ?

OpenStudy (anonymous):

for example if you call \(y=2^x\) then you have \[y^2-y-12=0\]

OpenStudy (anonymous):

you cannot take the log piece by piece here is a simple example we know \(3+7=10\) taking the log we have \[\log(3+7)=\log(10)\] not \[\log(3)+\log(7)=\log(10)\]

OpenStudy (christos):

that applies for nl(x) too?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[\log(a+b)\neq \log(a)+\log(b)\]

OpenStudy (christos):

ok please solve me this 2e^-x + 1 = 5

OpenStudy (christos):

I need to see something

OpenStudy (anonymous):

true of almost any function right? \[\sqrt{9+16}\neq \sqrt{9}+\sqrt{16}\]

OpenStudy (christos):

yes can you please solve me this, its an easy one, I just need to meisure a small logic that's behind

OpenStudy (anonymous):

\[2e^{-x+1}=5\]?

OpenStudy (anonymous):

or \[2e^{-x}+1=5\]?

OpenStudy (christos):

2e^(-x) + 1 = 5 sorry

OpenStudy (anonymous):

your first goal is to make it look like \[e^{-x}=\text{something}\] so you can solve by taking the log

OpenStudy (anonymous):

first step is to subtract 1 to get \[2e^{-x}=4\]

OpenStudy (anonymous):

divide by 2 and get \[e^{-x}=2\]

OpenStudy (anonymous):

in equivalent logarithmic for that means \[-x=\ln(2)\] and so \[x=-\ln(2)\]

OpenStudy (christos):

Then it goes x = ln(2^-1) and x = ln(1/2) :P

OpenStudy (anonymous):

if you like, yes they are the same number

OpenStudy (anonymous):

\[\log(\frac{1}{a})=-\log(a)\]

OpenStudy (anonymous):

however your first problem does not really involve logs

OpenStudy (christos):

log(a^-1)

OpenStudy (anonymous):

\[y^2-y-12=0\] \[(y-4)(y+3)=0\] \[y=4, y=-3\] \[2^x=4\] or \[2^x=-3\] the second one is not possible, and if \[2^x=4\] then \[x=2\]

OpenStudy (christos):

please give me the first step to solve (e^x - e^-x)/2 = 2